# Large Language Model-aided Edge Learning in Distribution System State Estimation

### v1

@@ II-A Preliminary of DSSE
where $V_{i},\theta_{i}$, represent the voltage magnitude and phase angle of bus $i$ respectively; $n$, and $m$ are the number of buses and measurements. Specifically, there is a measurement function $h(\cdot)$ that maps $\boldsymbol{x}$ to $\boldsymbol{z}$. In real systems, there is measurement noise, and thus the relationship between the measurement and states could be as follow:
$\boldsymbol{z}=h(\boldsymbol{x})+\boldsymbol{e},$ (2)
where $\boldsymbol{e}$ is the measurement error with zero means and a known variance $\boldsymbol{\sigma}$. Therefore, the state estimation problem is to find the function $h^{-1}(\boldsymbol{z})$ that maps the measurement to the state $\boldsymbol{x}$ while compensating the noise $\boldsymbol{e}$ In the distribution system, the measurement could be the different combination of the voltage magnitude or phase angle, active/reactive power injection, and active/reactive line power flows:
@@
where $P_{ij}$, $Q_{ij}$ represent the active power and reactive power flow from bus $i$ to bus $j$ respectively; $g_{ij}$, $b_{ij}$ are the line conductance and susceptance, which are the real and imaginary part of the line admittance $Y_{ij}$ respectively; $P_{inj,i},Q_{inj,i}$ are the active and reactive power injection of bus $i$.
@@ IV-C Performance of forecasting-aided state estimation
V Conclusion

### v2

@@ II-A Preliminary of DSSE
where $V_{i},\theta_{i}$, represent the voltage magnitude and phase angle of bus $i$ respectively; $n$, and $m$ are the number of buses and measurements. Specifically, there is a measurement function $h(\cdot)$ that maps $\boldsymbol{x}$ to $\boldsymbol{z}$. In the distribution system as depicted in 1, the measurement could be the different combination of the voltage magnitude or phase angle, active/reactive power injection, and active/reactive line power flows:
@@
where $P_{ij}$, $Q_{ij}$ represent the active power and reactive power flow from bus $i$ to bus $j$ respectively; $g_{ij}$, $b_{ij}$ are the line conductance and susceptance, which are the real and imaginary part of the line admittance $Y_{ij}$ respectively; $P_{inj,i},Q_{inj,i}$ are the active and reactive power injection of bus $i$. The first two subequations calculate the active and reactive power flowing from bus $i$ to bus $j$, based on the voltage levels, phase angle differences and electrical properties of the transmission line. And the last two subequations determines the net active power injected at bus $i$ by summing the differences between incoming and outgoing active power flows. They are fundamental to power flow analysis, which determines the voltage, current, and power flows in a power system under a given load and generation condition.
Figure 1: Power flow between buses.
@@ IV-C Performance of forecasting-aided state estimation
V Discussion
Although the proposed method can better handle the voltage magnitude and angle estimation tasks, and make an accurate state estimation with missing measurement, there are still several limitations.
First of all, the missing value prediction is implemented using the large language model, which could be cumbersome and require a lot of memory to be deployed, especially when using a complex model such as LLaMa 3.1, with a parameter of 405 billion. Therefore, the model quantization technique could be a potential future work for better application of large language models with limited hardware resources.
Secondly, only the missing values are considered in the real measurement, whereas there could be anomaly values or false data in the measurement. Therefore, integrating the anomaly detection function in the forecasting model could be more beneficial. A possible way is to consider the missing value imputing and anomaly detection a multitask learning problem, and use the uncertainty weight loss for training a comprehensive model to deal with missing and abnormal value simultaneously.
In addition, although the multitask learning improves the state estimation performance by finding the common features as well as the difference between tasks, it is still a data-driven approach. The black-box nature makes it hard to explain. Therefore, integrating physical information could benefit the data-driven method and provide more reliable results.
VI Conclusion